OpenStudy (anonymous):

Write a polynomial function of least degree with real coefficients in standard form with the given zeros: -1,4, square root of 3

OpenStudy (anonymous):

start with \[(x+1)(x-4)(x+\sqrt{3})(x-\sqrt{3})\] and then multiply out

OpenStudy (anonymous):

oh actually that is wrong. it says "real coefficients' not "integer coefficients"

OpenStudy (anonymous):

so if you take it literally you can use \[(x+1)(x-4)(x-\sqrt{3})\] it will be nice and ugly

OpenStudy (anonymous):

that's what i thought

OpenStudy (anonymous):

So would it be \[x ^3-3x ^{2}-4x-x ^{2}\sqrt{3}+3x \sqrt{3}+4\sqrt{3}\] or is that completely wrong?